show work please
4. Given a demand function:Q = 100 – 12p – 2p2, find the price elasticity of demand & op at p = 1.
Answer #1

Demand is Q = 100 - 12P - 2P^2

Derivative dQ/dP = -12 - 4P

Elasticity = dQ/dP * P/Q

When the price P = 1, Qd = 100 - 12*1 - 2*1^2 = 86 units

Hence elasticity of demand ed = (-12 - 4*1)*1/86

= -0.186

This is the value of elasticity of demand